OpenStudy (anonymous):

The rate of decay in the mass, M, of a radioactive substance is given by the differential equation dM /dt = -kM, where k is a positive constant. If the initial mass was 200g, then find the expression for the mass, M, at any time t.

OpenStudy (anonymous):

@Astrophysics @ganeshie8 @nincompoop

OpenStudy (anonymous):

\[\frac{ 1 }{ -k }\]\[\int\limits \frac{ dM }{ M } = \int\limits dt\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Yea you mean \[\frac{ -1 }{ k} \times \int\limits \frac{ dM }{ M}=\int\limits dt\] R Right?

OpenStudy (anonymous):

integrate both side and it is \[\frac{ -1 }{ k}( \ln M)=t\]

OpenStudy (anonymous):

okay how did you integrate the left side

OpenStudy (anonymous):

okay nvm i see what you did

OpenStudy (anonymous):

Cross multiply so you can get M by it self. \[-1(\ln M)=kt\]

OpenStudy (anonymous):

now i just solve for M right?

OpenStudy (anonymous):

ln M=-kt

OpenStudy (anonymous):

M = e ^-kt

OpenStudy (anonymous):

e both sides and it should look \[M=e ^{-kt}\]

OpenStudy (astrophysics):

Don't forget the constant you will need it for your initial amount

OpenStudy (anonymous):

Yea integrate both side and solve for M because he want to find an expression for M

OpenStudy (anonymous):

200 = e^-k(0) ...?

OpenStudy (anonymous):

Sorry had trouble with equation table.

OpenStudy (anonymous):

these are my answer choices by the way M = 200ln(kt) M = 2e−kt M = 200 ekt M = 200 e−kt

OpenStudy (anonymous):

is it D

OpenStudy (astrophysics):

yes

OpenStudy (anonymous):

Yes, because if it is M=200e^-kt

OpenStudy (anonymous):

okay thanks for the help guys :D

OpenStudy (anonymous):

if it is M=200e^-kt*

OpenStudy (anonymous):

you multiply it by 200 since it is decaying and it is the initial principal.

OpenStudy (astrophysics):

\[\frac{ dM }{ dt } = -kM \implies \int\limits \frac{ d M}{ M } = \int\limits - k dt \] \[\ln(M) = -k t+C\] \[\large e^{\ln(M)} = e^{-kt+C}\] \[M = e^{-kt}e^C\] let \[e^c = C\] so we have \[M = Ce^{-kt}\] after putting our initial conditions we get \[M = 200e^{-kt}\]

OpenStudy (astrophysics):

Just to note initial conditions is t = 0 meaning M = 200, so C = 200

OpenStudy (anonymous):

This method works too, but I just multiply the initial mass amount by the expression.

OpenStudy (anonymous):

Since he probably is taking a calculus class with some basic differential equation problems. From my experience, it is similar to when I took Calculus. Doubt this is a Differential Equation class questions. I never took DE.